OpenStudy (anonymous):

david has $17,000 to invest.he invests $12,000 in a mutual fund that pays 12% simple interest.if he wants to make at least $2200 in yearly interest at what minimum rate does the remainder of the money need to be invested

OpenStudy (tkhunny):

Solve for "Rate": 12000*0.12 + (17000-12000)*Rate = 2200 Look carefully at each piece. Make sure you know where it comes from.

OpenStudy (anonymous):

would that be 1440=5000+2200 what do i do then

OpenStudy (anonymous):

i mean 1440+5000=2200

OpenStudy (tkhunny):

?? Where did "Rate" go?

OpenStudy (anonymous):

im a lilttle confused what step should i take

OpenStudy (tkhunny):

You should solve for "Rate". It magically disappeared. 12000*0.12 + (17000-12000)*Rate = 2200 1440 + 5000*Rate = 2200 Now what?

OpenStudy (anonymous):

sorry im still lost i have no idea how to solve this problem

OpenStudy (tkhunny):

Background Check! What course are you in? Is it on-line? What section are you studying? Did you have to take Algebra before this course? I have to struggle with a few details if it appears you have been set up for failure by giving you a problem for which you have no reasonable chance for success.

OpenStudy (anonymous):

i am taking a college course in school elementary algebra this i am doing our midterm paper but this problem was not one we have done together so yes thats why im so confused this paper we were ablew to do for homework but is only used for extra points

OpenStudy (tkhunny):

Fair enough. Thank you for answering all those questions. $12000 earning 12% for one year yields $12000 * 12% = $12000 * 0.12 = $1440 The challenge there was converting from percent to decimal 12% ==> 0.12 In THIS problem, we have to odd items. There is a total of $17000 to INVEST. We already invested $12000. This leaves $5000 for the other investment. There is a total of $2200 to EARN. We already earned $1440. This leaves $760 for the other investment. The final decision is on this, then, what RATE do we need to get $5000 to earn $760? $5000 * Rate = $760 This is a division problem: Rate = $760 / $5000 = 0.152 = 15.2% All those words were wrapped up in that original equation. 12000*0.12 + (17000-12000)*Rate = 2200 1440 + 5000*Rate = 2200 5000*Rate = 2200-1440 = 760 Rate = 760/5000 = 0.152 = 15.2% The bad news is this. if you are having trouble solving this equation, you MAY need to back up and work a little harder on some algebra fundamentals. The good news is this, if this IS you working a little harder on your algebra, then you are totally awesome and keep up the good work!!

OpenStudy (anonymous):

a little of both im a freshman obviously been out of school about nine years so its taking some time to make it click back in our course its at a fast pace so i dont remember our teacher covering these problems thanks!!!1

OpenStudy (tkhunny):

Only 9? Then you are super awesome. Some leave it for 30 years and they totally regret it. Way to go getting back to it sooner. We'll help where we can. Keep trying. You'll get it!

OpenStudy (anonymous):

thanks